# Revenue Maximization Under Sequential Price Competition Via The Estimation Of $s$-Concave Demand Functions

### v3

@@ F.1 Proof of (1)
Then $\mathbf{y}^{\top}\mathbf{y}=y_{1}^{2}+\|\bm{\alpha}\|_{2}^{2}$. From the constraint in Equation˜27:
@@
where we defined the conditional center (the "shift")
@@ F.2 Proof of (2)
where we recall from Equation˜31 that $\mathbf{m}_{u}=\mathbf{m}+\Lambda\bm{\theta}\frac{\left(u-\bm{\theta}^{\top}\mathbf{m}\right)}{\bm{\theta}^{\top}\Lambda\bm{\theta}}$. Note that
@@
proving that $\psi_{i,\bm{\theta}}\in\mathrm{C}^{2}(\mathcal{U})$ for all $\bm{\theta}$. By ˜3.1 we have that $0<\underline{B}_{\psi_{i}}\leq\psi_{i}\leq\bar{B}_{\psi_{i}}$, which immediately implies that $0<\underline{B}_{\psi_{i}}\leq\psi_{i,\bm{\theta}}\leq\bar{B}_{\psi_{i}}$ uniformly in $\bm{\theta}$, proving point (2).
@@ F.4 Proof of (4)
where $c(\bm{\theta})\triangleq\frac{d}{du}\bm{\theta}_{i}^{\top}\mathbf{m}_{u}=\frac{\bm{\theta}_{i}^{\top}\Lambda\bm{\theta}}{\bm{\theta}^{\top}\Lambda\bm{\theta}}$ as defined in (33). Then we only need to prove that $\bm{\theta}_{i}^{\top}\Lambda\bm{\theta}>0$ for $\bm{\theta}$ sufficiently close to $\bm{\theta}_{i}$. Recall by ˜5.1 that $\Lambda\in\mathbb{R}^{N\times N}$ is symmetric positive definite matrix with $c_{\min}\mathbb{I}\preccurlyeq\Lambda\preccurlyeq c_{\max}\mathbb{I}$ for some $0<c_{\min}\leq c_{\max}<\infty$, and that $\|\bm{\theta}\|_{2}=\|\bm{\theta}_{i}\|_{2}=1$. We are interested in the conditions under which
@@ F.5 Proof of (5)
is convex on $(0,\infty)\times(0,\infty)$, and then we can apply Niculescu & Persson (2006, Theorem 3.5.3), which is an extension of the Jensen’s inequality for in $2$ variables. ∎
Appendix G Proof of Theorem˜5.4
@@
where we used ˜3.2: $0<\underline{B}_{\psi_{i}^{\prime}}\leq\psi_{i}^{\prime}\leq\bar{B}_{\psi_{i}^{\prime}}$ and $|\psi_{i}^{\prime\prime}|\leq B_{\psi_{i}^{\prime\prime}}$ on $\mathcal{U}$ for some (unknown) $\underline{B}_{\psi_{i}^{\prime}},\bar{B}_{\psi_{i}^{\prime}},B_{\psi_{i}^{\prime\prime}}>0$. This implies
@@
If Assumptions in the Theorem hold, for $T$ sufficiently large we have $\mathbb{E}[\sup_{\mathbf{p}\in\mathcal{P}}|\psi_{i}(\bm{\theta}_{i}^{\top}\mathbf{p})-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(\widetilde{\bm{\theta}}_{i}^{\top}\mathbf{p})|]\lesssim\sqrt{N}(\tfrac{\log\tau}{\tau})^{2/5}$. Moreover, there exists a unique value $\kappa_{i}^{\star}\in(0,1)$ that minimizes $\mathbb{E}[\sup_{\mathbf{p}\in\mathcal{P}}|\psi_{i}(\bm{\theta}_{i}^{\top}\mathbf{p})-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(\widetilde{\bm{\theta}}_{i}^{\top}\mathbf{p})|]$. This value satisfies the implicit equation:
@@
By Lemma˜G.1, Lemma˜G.2 have that for $t\geq\tau+1$ with $\tau$ sufficiently large,
@@
$\lesssim\tau+(T-\tau)N^{3/2}\left(\frac{\log\tau}{\tau}\right)^{2/5}+\sum_{t=\tau+1}^{T}L^{2(t-\tau-1)}_{\bm{\Gamma}}$
$\lesssim\tau+TN^{3/2}\left(\frac{\log\tau}{\tau}\right)^{2/5}+\sum_{j=0}^{T-\tau-1}L^{2j}_{\bm{\Gamma}}$
$=T^{\xi}+TN^{3/2}\left(\frac{\log T^{\xi}}{T^{\xi}}\right)^{2/5}+\frac{1-L^{2(T-\tau)}_{\bm{\Gamma}}}{1-L^{2}_{\bm{\Gamma}}}$
$\leq T^{\xi}+T^{1-2\xi/5}N^{3/2}\left(\log T\right)^{2/5}+1.$
Then we get
$\operatorname{Reg}_{i}(T)\lesssim T^{\xi}+T^{1-\nicefrac{{2\xi}}{{5}}}N^{\nicefrac{{3}}{{2}}}(\log T)^{\nicefrac{{2}}{{5}}}.$
As a consequence, substituting back the value $\tau\propto T^{\xi}$ into Appendix˜G, we get
$\mathbb{E}[\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}]$ $\lesssim N^{\nicefrac{{3}}{{2}}}(\tfrac{\log\tau}{\tau})^{\nicefrac{{2}}{{5}}}+L^{2(t-\tau-1)}_{\bm{\Gamma}}$
$\lesssim N^{\nicefrac{{3}}{{2}}}(\tfrac{\log T^{\xi}}{T^{\xi}})^{\nicefrac{{2}}{{5}}}+L^{2(t-T^{\xi}-1)}_{\bm{\Gamma}}$
$=N^{\nicefrac{{3}}{{2}}}T^{-\nicefrac{{2\xi}}{{5}}}(\log T)^{\nicefrac{{2}}{{5}}}+L^{2(t-T^{\xi}-1)}_{\bm{\Gamma}}$,
@@
$\mathbb{E}[\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}]\lesssim N^{\nicefrac{{3}}{{2}}}T^{-\nicefrac{{2\xi}}{{5}}}(\log T)^{\nicefrac{{2}}{{5}}}+L^{2(T-T^{\xi}-1)}_{\bm{\Gamma}}$.
This complete the proof.
G.1 Proof of Lemma˜G.1
Let $\tau$ be the minimum exploration phase among the $N$ firms and $t\in\{\tau+1,\tau+2,\dots,T\}$. Let $n_{i}^{(1)}$ be the length of the exploration phase to estimate $\bm{\theta}_{i}$ and $n_{i}^{(2)}$ the remaining length for estimating the link function $\psi_{i}$. Define the event $\mathcal{E}_{i}=\{\|\widetilde{\bm{\theta}}_{i}-\bm{\theta}_{i}\|\leq R_{n_{i}^{(1)}}\}$ where $R_{n_{i}^{(1)}}=\mathscr{B}_{i}\sqrt{\frac{N\log n_{i}^{(1)}}{n_{i}^{(1)}}}$, where $\mathscr{B}_{i}$ is defined in Proposition˜D.2, and
@@
By Proposition˜D.2 we have $\mathbb{E}[\mathscr{R}_{i}(p_{i}|\mathbf{p}_{-i})\mathbb{I}(\mathcal{E}_{i}^{c})]\lesssim\mathbb{P}(\mathcal{E}_{i}^{c})=Q_{n_{i}^{(1)}}=2e^{-c_{1}\varsigma_{\min}^{2}n_{i}^{(1)}/16}+\frac{2}{n_{i}^{(1)}}$.
@@
Analyzing $A$ on $\mathcal{E}_{i}$: Define the event $\mathcal{Q}_{i}=\{\sup_{u\in\mathcal{U}}|\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(u)-\psi_{i,\widetilde{\bm{\theta}}_{i}}(u)|\leq\mathscr{C}_{i}\rho_{n_{i}^{(2)}}^{2/5}\}$. For $n_{i}^{(2)}$ sufficiently large, by Theorem˜D.5 we have that for every $\gamma>4$
@@
Analyzing $B$ on $\mathcal{E}_{i}$: By Proposition˜D.3, $\psi_{i,\widetilde{\bm{\theta}}_{i}}$ is Lipschitz, then
@@
Analyzing $C$ on $\mathcal{E}_{i}$: By Proposition˜D.3 we have
@@
$\mathscr{R}_{i}(p_{i}|\mathbf{p}_{-i})\lesssim\left(\frac{\log(\tau(1-\kappa_{i}))}{\tau(1-\kappa_{i})}\right)^{2/5}+\left(N\frac{\log(\tau\kappa_{i})}{\tau\kappa_{i}}\right)^{1/2}\lesssim\sqrt{N}\left(\frac{\log n_{i}^{(2)}}{n_{i}^{(2)}}\right)^{2/5}.$
G.2 Proof of Lemma˜G.2
@@
which is strictly positive by ˜3.2. Then we have

### v4

@@ F.1 Proof of (1)
Then $\mathbf{y}^{\top}\mathbf{y}=y_{1}^{2}+\|\bm{\alpha}\|_{2}^{2}$. From the constraint in Equation 27:
@@
where we defined the conditional center (the ”shift”)
@@ F.2 Proof of (2)
where we recall from Equation 31 that $\mathbf{m}_{u}=\mathbf{m}+\Lambda\bm{\theta}\frac{\left(u-\bm{\theta}^{\top}\mathbf{m}\right)}{\bm{\theta}^{\top}\Lambda\bm{\theta}}$. Note that
@@
proving that $\psi_{i,\bm{\theta}}\in\mathrm{C}^{2}(\mathcal{U})$ for all $\bm{\theta}$. By 3.1 we have that $0<\underline{B}_{\psi_{i}}\leq\psi_{i}\leq\bar{B}_{\psi_{i}}$, which immediately implies that $0<\underline{B}_{\psi_{i}}\leq\psi_{i,\bm{\theta}}\leq\bar{B}_{\psi_{i}}$ uniformly in $\bm{\theta}$, proving point (2).
@@ F.4 Proof of (4)
where $c(\bm{\theta})\triangleq\frac{d}{du}\bm{\theta}_{i}^{\top}\mathbf{m}_{u}=\frac{\bm{\theta}_{i}^{\top}\Lambda\bm{\theta}}{\bm{\theta}^{\top}\Lambda\bm{\theta}}$ as defined in (33). Then we only need to prove that $\bm{\theta}_{i}^{\top}\Lambda\bm{\theta}>0$ for $\bm{\theta}$ sufficiently close to $\bm{\theta}_{i}$. Recall by 5.1 that $\Lambda\in\mathbb{R}^{N\times N}$ is symmetric positive definite matrix with $c_{\min}\mathbb{I}\preccurlyeq\Lambda\preccurlyeq c_{\max}\mathbb{I}$ for some $0<c_{\min}\leq c_{\max}<\infty$, and that $\|\bm{\theta}\|_{2}=\|\bm{\theta}_{i}\|_{2}=1$. We are interested in the conditions under which
@@ F.5 Proof of (5)
is convex on $(0,\infty)\times(0,\infty)$, and then we can apply niculescu2006convex, which is an extension of the Jensen’s inequality for in $2$ variables. ∎
Appendix G Proof of Theorem 5.4
@@
where we used 3.2: $0<\underline{B}_{\psi_{i}^{\prime}}\leq\psi_{i}^{\prime}\leq\bar{B}_{\psi_{i}^{\prime}}$ and $|\psi_{i}^{\prime\prime}|\leq B_{\psi_{i}^{\prime\prime}}$ on $\mathcal{U}$ for some (unknown) $\underline{B}_{\psi_{i}^{\prime}},\bar{B}_{\psi_{i}^{\prime}},B_{\psi_{i}^{\prime\prime}}>0$. This implies
@@
If Assumptions in the Theorem hold, for $T$ sufficiently large we have $\mathbb{E}[\sup_{\mathbf{p}\in\mathcal{P}}|\psi_{i}(\bm{\theta}_{i}^{\top}\mathbf{p})-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(\widetilde{\bm{\theta}}_{i}^{\top}\mathbf{p})|]\lesssim\mathscr{X}_{i}\sqrt{N}(\tfrac{\log\tau}{\tau})^{2/5}$, where $\mathscr{X}_{i}=\max\left\{\mathscr{C}_{i},\mathscr{B}_{i}\right\}$, and $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ are defined in (17) and (21), respectively. Moreover, there exists a unique value $\kappa_{i}^{\star}\in(0,1)$ that minimizes $\mathbb{E}[\sup_{\mathbf{p}\in\mathcal{P}}|\psi_{i}(\bm{\theta}_{i}^{\top}\mathbf{p})-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(\widetilde{\bm{\theta}}_{i}^{\top}\mathbf{p})|]$. This value satisfies the implicit equation:
@@
By Lemma G.1, Lemma G.2 have that for $t\geq\tau+1$ with $\tau$ sufficiently large,
@@
$\lesssim\tau+(T-\tau)\mathscr{X}_{i}N^{3/2}\left(\frac{\log\tau}{\tau}\right)^{2/5}+\sum_{t=\tau+1}^{T}L^{2(t-\tau-1)}_{\bm{\Gamma}}$
$\lesssim\tau+T\mathscr{X}_{i}N^{3/2}\left(\frac{\log\tau}{\tau}\right)^{2/5}+\sum_{j=0}^{T-\tau-1}L^{2j}_{\bm{\Gamma}}$
$=T^{\xi}+T\mathscr{X}_{i}N^{3/2}\left(\frac{\log T^{\xi}}{T^{\xi}}\right)^{2/5}+\frac{1-L^{2(T-\tau)}_{\bm{\Gamma}}}{1-L^{2}_{\bm{\Gamma}}}$
$\leq T^{\xi}+\mathscr{X}_{i}T^{1-2\xi/5}N^{3/2}\left(\log T\right)^{2/5}+1.$
Note that we are retaining the dependence on $\mathscr{X}_{i}=\mathscr{X}_{i}(s_{i})$ to demonstrate the way that $s_{i}$ affects the rate of the regret. Substituting back the value $\tau\propto T^{\xi}$ into Appendix G, we get
$\mathbb{E}[\|\mathbf{p}^{(t)}-\mathbf{p}^{*}\|_{2}^{2}]$ $\lesssim\mathscr{X}_{i}N^{\nicefrac{{3}}{{2}}}(\tfrac{\log\tau}{\tau})^{\nicefrac{{2}}{{5}}}+L^{2(t-\tau-1)}_{\bm{\Gamma}}$
$\lesssim\mathscr{X}_{i}N^{\nicefrac{{3}}{{2}}}(\tfrac{\log T^{\xi}}{T^{\xi}})^{\nicefrac{{2}}{{5}}}+L^{2(t-T^{\xi}-1)}_{\bm{\Gamma}}$
$=\mathscr{X}_{i}N^{\nicefrac{{3}}{{2}}}T^{-\nicefrac{{2\xi}}{{5}}}(\log T)^{\nicefrac{{2}}{{5}}}+L^{2(t-T^{\xi}-1)}_{\bm{\Gamma}}$,
@@
$\mathbb{E}[\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}]\lesssim\mathscr{X}_{i}N^{\nicefrac{{3}}{{2}}}T^{-\nicefrac{{2\xi}}{{5}}}(\log T)^{\nicefrac{{2}}{{5}}}+L^{2(T-T^{\xi}-1)}_{\bm{\Gamma}}$.
This completes the proof.
Remark G.3 (Dependence of the regret and NE convergence rate on $s_{i}$).
The value of $\xi$ that minimizes the expected regret $\operatorname{Reg}_{i}(T)$ is obtained by equalizing the exponents of $T$ in the upper bound $T^{\xi}+T^{1-\nicefrac{{2\xi}}{{5}}}$, ignoring logarithmic factors. Solving this balance gives $\xi=5/7$. Consequently, for each $i\in\mathcal{N}$ we obtain
$\operatorname{Reg}_{i}(T)$ $=O\!\left(\mathscr{X}_{i}\,N^{\nicefrac{{3}}{{2}}}\,T^{\nicefrac{{5}}{{7}}}(\log T)^{\nicefrac{{2}}{{5}}}\right),$
$\mathbb{E}\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}$ $=O\!\left(\mathscr{X}_{i}\,N^{\nicefrac{{3}}{{2}}}\,T^{-\nicefrac{{2}}{{7}}}(\log T)^{\nicefrac{{2}}{{5}}}+L_{\bm{\Gamma}}^{\,2T+o(T)}\right),\qquad T\to\infty.$
These bounds retain the constants $\mathscr{X}_{i}$ and $L_{\bm{\Gamma}}$ because they depend on the parameters $s_{1},\dots,s_{N}$ (or equivalently $c_{1},\dots,c_{N}$). Note that the dependence on $\{s_{j}\}_{j\in[N]}$ enters only through the multiplicative constant $\mathscr{X}_{i}$ and the additive constant $L_{\bm{\Gamma}}$, never through the powers of $T$. This makes explicit how the choice of $s_{i}$ influences the regret and the convergence to the NE. Below we show that if $s_{i}$ are the optimal-concave parameters (as defined in Definition L.3), i.e.
$s_{i}=\sup\Bigl\{t_{i}:\sup_{u\in\mathcal{U}_{i}}\bigl[\psi_{i}(u)\psi_{i}^{\prime\prime}(u)+(t_{i}-1)(\psi_{i}^{\prime}(u))^{2}\bigr]\leq 0\Bigr\},$
then any misspecification $s_{i}^{\prime}\leq s_{i}$ satisfying
$\sup_{i\in\mathcal{N}}\frac{|1-\nicefrac{{1}}{{s_{i}^{\prime}+1}}|\;\|\bm{\gamma}_{i}\|_{1}}{\beta_{i}}<1$
does not affect the regret or the NE convergence rates.
  1. (1)
    Multiplicative constants $\mathscr{X}_{i}$. The constant $\mathscr{X}_{i}=\{\mathscr{C}_{i},\mathscr{B}_{i}\}$ contains $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ (defined in (17)–(21)), which are the multiplicative constants in the concentration bounds for
    $\|\psi_{i}-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}\|_{\infty}\qquad\text{and}\qquad\|\bm{\theta}_{i}-\widetilde{\bm{\theta}}_{i}\|_{2},$
    respectively. Any value $s_{i}^{\prime}\neq s_{i}$ does not affect the rate of linear estimator $\widetilde{\bm{\theta}}_{i}$, but it could introduce a bias in the estimator $\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}$. Precisely, if we choose $s_{i}^{\prime}\leq s_{i}$, the inclusion property of Proposition L.1 implies that an $s_{i}$-optimal-concave function remains $s_{i}^{\prime}$-concave. Thus no bias is introduced in the concentration of $\|\psi_{i}-\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}\|_{\infty}$, whereas a choice $s_{i}^{\prime}>s_{i}$ would. Then for $s_{i}^{\prime}\leq s_{i}$, since $\mathscr{B}_{i}$ and $\mathscr{C}_{i}$ are uniformly bounded in $s_{i}$, i.e. $\mathscr{B}_{i},\mathscr{C}_{i}<C$ for some $C>0$ independent of $s_{i}$, they do not affect the regret or NE rates.
  2. (2)
    Additive constant $L_{\bm{\Gamma}}$. The constant $L_{\bm{\Gamma}}$ is the contraction modulus of the best-response map and determines the geometric decay term in the NE convergence. Convergence holds whenever $L_{\bm{\Gamma}}\in(0,1)$, or equivalently $\sup_{i\in\mathcal{N}}\nicefrac{{\|g_{i}^{\prime}\|_{\infty}\,\|\bm{\gamma}_{i}\|_{1}}}{{\beta_{i}}}<1.$ Because $g_{i}^{\prime}(u)=1-\nicefrac{{1}}{{\varphi_{i}^{\prime}(\varphi_{i}^{-1}(u))}}$, we have
    $1-\frac{1}{c_{i}}<g_{i}^{\prime}(u)<1,\qquad c_{i}=s_{i}+1.$
    Thus any $s_{i}^{\prime}$ satisfying $\sup_{i\in\mathcal{N}}\nicefrac{{|1-\nicefrac{{1}}{{s_{i}^{\prime}+1}}|\;\|\bm{\gamma}_{i}\|_{1}}}{{\beta_{i}}}<1$ preserves contraction. If $L_{\bm{\Gamma}}\geq 1$, contraction fails and the regret guarantees break down.
Combining both components, we conclude that as long as the sellers choose $s_{i}^{\prime}\leq s_{i}$ (where $s_{i}$ are the optimal-concave parameters of $\psi_{i}$) and $\sup_{i\in\mathcal{N}}\nicefrac{{|1-\nicefrac{{1}}{{s_{i}^{\prime}+1}}|\;\|\bm{\gamma}_{i}\|_{1}}}{{\beta_{i}}}<1$, the convergence rates of both the regret and the NE remain unchanged.
G.1 Proof of Lemma G.1
Let $\tau$ be the minimum exploration phase among the $N$ firms and $t\in\{\tau+1,\tau+2,\dots,T\}$. Let $n_{i}^{(1)}$ be the length of the exploration phase to estimate $\bm{\theta}_{i}$ and $n_{i}^{(2)}$ the remaining length for estimating the link function $\psi_{i}$. Define the event $\mathcal{E}_{i}=\{\|\widetilde{\bm{\theta}}_{i}-\bm{\theta}_{i}\|\leq R_{n_{i}^{(1)}}\}$ where $R_{n_{i}^{(1)}}=\mathscr{B}_{i}\sqrt{\frac{N\log n_{i}^{(1)}}{n_{i}^{(1)}}}$, where $\mathscr{B}_{i}$ is defined in Proposition D.3, and
@@
By Proposition D.3 we have $\mathbb{E}[\mathscr{R}_{i}(p_{i}|\mathbf{p}_{-i})\mathbb{I}(\mathcal{E}_{i}^{c})]\lesssim\mathbb{P}(\mathcal{E}_{i}^{c})=Q_{n_{i}^{(1)}}=2e^{-c_{1}\varsigma_{\min}^{2}n_{i}^{(1)}/16}+\frac{2}{n_{i}^{(1)}}$.
@@
Analyzing $A$ on $\mathcal{E}_{i}$: Define the event $\mathcal{Q}_{i}=\{\sup_{u\in\mathcal{U}}|\widehat{\psi}_{i,\widetilde{\bm{\theta}}_{i}}(u)-\psi_{i,\widetilde{\bm{\theta}}_{i}}(u)|\leq\mathscr{C}_{i}\rho_{n_{i}^{(2)}}^{2/5}\}$. For $n_{i}^{(2)}$ sufficiently large, by Theorem D.6 we have that for every $\gamma>4$
@@
Analyzing $B$ on $\mathcal{E}_{i}$: By Proposition D.4, $\psi_{i,\widetilde{\bm{\theta}}_{i}}$ is Lipschitz, then
@@
Analyzing $C$ on $\mathcal{E}_{i}$: By Proposition D.4 we have
@@
$\mathscr{R}_{i}(p_{i}|\mathbf{p}_{-i})\lesssim\mathscr{C}_{i}\left(\frac{\log(\tau(1-\kappa_{i}))}{\tau(1-\kappa_{i})}\right)^{2/5}+\mathscr{B}_{i}\left(N\frac{\log(\tau\kappa_{i})}{\tau\kappa_{i}}\right)^{1/2}\lesssim\mathscr{X}_{i}\sqrt{N}\left(\frac{\log n_{i}^{(2)}}{n_{i}^{(2)}}\right)^{2/5}.$
where $\mathscr{X}_{i}=\max\left\{\mathscr{C}_{i},\mathscr{B}_{i}\right\}$.
G.2 Proof of Lemma G.2
@@
which is strictly positive by 3.2. Then we have